# Context versus Prior Knowledge in Language Models

### v1

@@ Abstract
To answer a question, language models often need to integrate prior knowledge learned during pretraining and new information presented in context. We hypothesize that models perform this integration in a predictable way across different questions and contexts: models will rely more on prior knowledge for questions about entities (e.g., persons, places, etc.) that they are more familiar with due to higher exposure in the training corpus, and be more easily persuaded by some contexts than others. To formalize this problem, we propose two mutual information-based metrics to measure a model's dependency on a context and on its prior about an entity: first, the persuasion score of a given context represents how much a model depends on the context in its decision, and second, the susceptibility score of a given entity represents how much the model can be swayed away from its original answer distribution about an entity. Following well-established measurement modeling methods, we empirically test for the validity and reliability of these metrics. Finally, we explore and find a relationship between the scores and the model's expected familiarity with an entity, and provide two use cases to illustrate their benefits.
@@ 1 Introduction
Language models have displayed remarkable abilities to answer factual queries about entities, suggesting that they encode knowledge about these entities learned during pretraining (Petroni et al., 2019; Brown et al., 2020; Roberts et al., 2020; Geva et al., 2021). For prompts that extend a question with additional information or context, the model can draw on both its prior knowledge and the additional context in order to answer the query (Kwiatkowski et al., 2019; Joshi et al., 2017; Berant et al., 2013; Kasai et al., 2023). While previous research has investigated how often a model will rely on prior knowledge over conflicting contextual information in answering questions (Longpre et al., 2021), we hypothesize that models will not behave identically for all contexts and entities. For example, if a language model is prompted with Harry hugged Voldemort. How friendly are Harry Potter and Lord Voldemort?, we might expect the prior knowledge learned from training data describing the rivalry between these two characters to significantly influence the model's answer. However, if the model lacks a strong prior on, say, Susie and Alia, then we might expect its answer to be primarily context-driven when prompted with Susie hugged Alia. How friendly are Susie and Alia?.
@@
We formalize this problem through the lens of evaluating the change in a model's answer distribution for different contexts and entities. We present two mutual information-based metrics that allow us to explore differences in the effect of specific contexts on model behavior for different entities. The persuasion score of a given context measures how much a model's answer distribution is affected by the context when prompted with a particular query about a given entity. The susceptibility score of a given entity measures how much the model's answer distribution can be swayed for a particular query about that entity, marginalized over all contexts. Given their basis on mutual information, these metrics are both faithful and comprehensive reflections of model behavior. We offer empirical evidence of the validity and reliability of these metrics by comparing them against similar metrics and showing their robustness (i.e., low variance) to paraphrases and different samples.
@@
Using these metrics and a synthetic dataset111https://github.com/kdu4108/measureLM covering 122 topic areas extracted from the YAGO knowledge graph (Suchanek et al., 2007), we study the effect of a variety of factors on the behavior of pretrained models (ranging from 70m to 12b parameters) across contexts and entities. We find evidence that relevant contexts are consistently more persuasive than irrelevant ones, and assertive contexts are more persuasive than less assertive ones for yes/no questions. In a deep dive for one model, we provide evidence that entities appearing frequently in the training data, and thus that we expect the model to be familiar with, have lower susceptibility scores. Since calculating more complex statistics (i.e., relation-specific entity co-occurrence) from noisy training data can be challenging, we further support our findings by obtaining the same result with a relation-dependent measure of expected familiarity based on the entity's degree in the YAGO knowledge graph.
@@
We further conduct case studies to demonstrate how these metrics could be useful in an applied setting. In a study on friend–enemy stance measurement, we find evidence that enemy duos are less susceptible than friend duos. Applying our metrics to gender bias, we find evidence for a difference in susceptibility between stereotypically masculine and feminine names with respect to gender-biased contexts. Through this, we demonstrate how our proposed metrics can be used to better analyze the effects of context and prior knowledge, with the potential to apply this knowledge to obtain greater control over model behavior.
@@ 2.1 Entity Bias
Loosely defined across studies as model bias from spurious correlations between entity mentions and a target characteristic in the training data, entity bias has been mainly examined in relation extraction (RE), where a model extracts relationships between entities from text (Zelenko et al., 2002; Zhou et al., 2005). Different studies on RE have attempted to either mitigate (Zhang et al., 2017, 2018; Peng et al., 2020; Wang et al., 2022) or leverage (Yamada et al., 2020; Zhou and Chen, 2022) entity bias for improved performance on the task. Wang et al. (2022) note that entities can carry both useful semantic information about their roles (e.g., whether the entity is a person or a place) and spurious information which can bias the model toward relations not mentioned in the target sentence (e.g., the model may link Switzerland with the relation countries_of_residence at inference-time for a sentence that does not mention this relation, due to their frequent co-occurrence in the training data). Additional studies on machine reading comprehension (MRC) tasks, e.g., SQuAD (Rajpurkar et al., 2016), TriviaQA (Joshi et al., 2017), and NaturalQuestions (Kwiatkowski et al., 2019), have found substituting different entity names can result in meaningful changes in model predictions and overall evaluation performance (Yan et al., 2022).
@@ 2.2 Context and Entity Bias
The existence of entity bias naturally raises the question of how it interacts with context to shape a model's response. Longpre et al. (2021) approach this by inducing and exploring knowledge conflicts, where a context preceding a query proposes information that conflicts with a model's prior knowledge about that query. They measure the model's reliance on pretrained entity bias and context by computing the memorization ratio: the proportion of knowledge-conflict examples for which the model maintains its answer from prior knowledge. Pezeshkpour (2023) measures a model's prior knowledge of a fact by comparing the entropy of a queried model's answer distribution before and after stating the fact in context. Several studies further propose interventions to reduce entity bias and favor in-context information. These approaches include prompting (Zhou et al., 2023; Onoe et al., 2023), modifying training data (Wang et al., 2023), fine-tuning (Li et al., 2023), and neuron-level interventions during inference (Yu et al., 2023).
However, the metrics used to quantify model reliance on context and entity bias in these works (excepting (Pezeshkpour, 2023)—memorization ratio and empirical task performance on existing RE or MRC tasks—are limited in several ways. First, these works do not find entity-specific or context-specific metrics for how strong the entity bias or context persuasiveness is. Instead, these metrics produce only a single number which summarizes the model's overall reliance on entity bias. Second, these metrics can be unreliable with high variance since they sample from the model's output distribution. Third, their setups are limited to adversarial cases in which a context disagrees with the entity bias. We may well wish to measure the interplay of context and entity bias in other cases, such as when context reinforces entity bias or when they do not clearly disagree. Therefore, we seek to define a metric that measures how much the model depends on a given context or entity with rigorous, theoretically grounded interpretations.
@@ 3 Our Formalization
We now formalize the problem setting in which we define metrics for the susceptibility of an entity and the persuasiveness of a context for a given model. Let be an alphabet. Consider a language model over , i.e., is a distribution over . Furthermore, we assume that was estimated from a corpus . Let that corresponds to string representations of entities. Let be a set of slotted query templates of type that fill the slot with the argument to each . We can think of a query template as slotting an entity into a query, e.g., slotting the entity Slovenia into the query template The capital of is produces the string The capital of Slovenia is.222In practice, an entity can have different verbalizations, e.g., Sherlock and Sherlock Holmes. Our formalization does not group different verbalizations into the same entity, which might be useful to explore in future work.
Now, consider three random variables. First, let be a -valued random variable that stands for a context. Second, let be a -valued random variable whose values are answers. Let be a -valued random variable. The pushforward , then, is a -valued random variable over slottings of entities according to query . The random variables , and are jointly distributed according to the following probability
where is string concatenation. To formalize persuasion and susceptibility, we make use of the joint distribution heavily in the subsequent subsections.
@@ 3.1 Persuasion Score
Formally, for each context prepended to a query template to an entity , we wish to assign a persuasion score to represent how successful that context is at altering a model's answer distribution. This score depends on the specific queried entity , because contexts themselves are often entity-dependent. Intuitively, a context's persuasion score should measure how much the log probability distribution of possible answers changes, averaged across all possible answers. More precisely, we define our persuasion score as the half-pointwise mutual information (half-PMI) between the context and the answer random variable , conditioned on the fixed query about an entity:
where we define the following marginal distribution
The persuasion score of a context can then be interpreted as the degree (in nats) to which the context was able to change the model's answer distribution when prepended to a query. When the persuasion score is at its lower bound of 0 nats, it indicates the context is completely unpersuasive, i.e., it did not change the model's answer distribution at all. Contexts with higher persuasion scores are more persuasive, i.e., they change the answer distribution more, which is consistent when viewed through the lens of KL-divergence.333We prove further equivalencies of half-PMI and other concepts in information theory in App. A. Due to its information-theoretic foundation, faithfulness to the actual model distribution, and interpretable units, when framed in existing measurement modeling methods, we argue this metric exhibits construct validity via face validity, i.e., the persuasion score is a plausible and meaningful measure for the persuasiveness of a context (Loevinger, 1957; Messick, 1987; Jackman, 2008; Hand, 2004; Quinn et al., 2010; Jacobs and Wallach, 2021).
@@ 3.2 Susceptibility Score
For a given query template applied to an entity , we further wish to assign a susceptibility score to which represents how easy it is to change a model's answer distribution. Intuitively, the susceptibility score should measure how much the log probability of an answer, when prompted with just the query, differs from the log probability of that answer when prompted with additional context, averaged across all possible contexts and answers. More precisely, we define the susceptibility score as the mutual information between the context and answer random variables, conditioned on a fixed query about an entity:
where is the marginal distribution over contexts. Equivalent to , the susceptibility score represents the reduction in answer distribution uncertainty (in nats) when a query is preceded by a context, and in a well-grounded statistical sense represents the dependence of answer on context , conditioned on the query–entity pair . A high susceptibility score means the model is highly influenced by context for the query about that entity, with its upper bound of indicating that context fully determines the answer. A low score signals the model's response is robust to context, with its lower bound of indicating no influence of context on the answer distribution. This information-theoretic metric enables us to answer the question ``how much does the model's answer depend on its entity bias?'' with an information-theoretically grounded, interpretable scale based on the model's full behavior. These qualities support our claim that susceptibility scores are meaningful and plausible and, thus, under measurement modeling frameworks, exhibit construct validity and face validity.
@@ 4.1 Setup
For each of the 122 different relations from the YAGO knowledge graph, we collect 100 entities44450 are real entities (e.g., Adele) sampled from YAGO, and 50 are fake entities (e.g., Udo König) of the same entity class (e.g., Person) generated with GPT-4 (OpenAI, 2023). and construct 600 random contexts from relation-specific context templates such that each entity is mentioned in 6 contexts. The 600 contexts are evenly distributed between assertive,555E.g., Definitely, the capital of {entity} is {answer}. base,666E.g., The capital of {entity} is {answer}. and negation777E.g., The capital of {entity} is not {answer}. context types. We construct four query forms of each relation: two closed questions (i.e., yes–no) and two open questions.
Using these samples, we compute persuasion scores for each context according to Eq. 2 and susceptibility scores for each entity according to Eq. 4 for each of the four query forms for six Pythia models of different sizes88870m, 410m, 1.4b, 2.8b, 6.9b, 12b (8-bit quantized) trained on the deduplicated Pile (Wolf et al., 2020; Dettmers et al., 2022; Biderman et al., 2023). Due to computational constraints, we approximate the model's answer distribution with the next token distribution over the model's vocabulary.999While it would be more precise to estimate the answer distribution by repeatedly sampling many model outputs, this is very computationally expensive. The detailed setup can be found in App. C. In the following sections, we describe further analyses and comparisons of these scores to validate the metrics and demonstrate how they can characterize model behavior.
@@ 4.2.1 Validating Persuasion Scores
According to existing measurement modeling methods, observing a relationship between a new metric and existing ones would serve as additional evidence that the metric is meaningful. To this end, we explore the relationship between persuasion scores and whether the model outputted an answer consistent with the context. We expect that contexts with higher persuasion scores should also be more successful at convincing the model to answer in agreement with the context. To test this, we use the Pythia-6.9b-deduped model to generate a model answer for each prepended context to a query–entity pair from § 4.1. We map each answer to whether it agrees with the context, the original answer, or neither via simple string matching. We then run permutation tests () with the Benjamini-Hochberg (BH) correction (Benjamini and Hochberg, 1995) for the questions in each of the 122 query topics to test for whether contexts eliciting context-agreeing answers have higher persuasion scores than contexts which fail to change the answer from the original. Within a query topic, we run separate tests for open and closed queries, since the entropy of the answer distribution for closed queries tends to be much lower than the entropy for open queries.
We find that for of open queries, contexts that persuade the model to output the in-context answer have significantly higher persuasion scores than non-persuasive ones. Curiously, this behavior holds for only of closed queries. We discuss this surprising result more in § 6. We further show a summary of the mean persuasion scores for the different kinds of elicited answers in Fig. 2.
@@
To show construct reliability, we provide evidence that persuasion scores have low variance across different samples of contexts and different phrasings of the same query (i.e., query forms) for the Pythia-6.9b-deduped model. Since computing the persuasion score of a context requires marginalizing over all contexts (Eq. 3), we want to show that this metric is robust to different contexts sampled from the same distribution. The metric should also be robust to different phrasings of the same question, e.g., the persuasion score of The capital of Slovenia is Gotham City. should be similar when prepended to either Is Slovenia's capital Ljubljana? or Is Ljubljana the capital of Slovenia?.

### v3

@@ Abstract
To answer a question, language models often need to integrate prior knowledge learned during pretraining and new information presented in context. We hypothesize that models perform this integration in a predictable way across different questions and contexts: models will rely more on prior knowledge for questions about entities (e.g., persons, places, etc.) that they are more familiar with due to higher exposure in the training corpus, and be more easily persuaded by some contexts than others. To formalize this problem, we propose two mutual information-based metrics to measure a model’s dependency on a context and on its prior about an entity: first, the persuasion score of a given context represents how much a model depends on the context in its decision, and second, the susceptibility score of a given entity represents how much the model can be swayed away from its original answer distribution about an entity. We empirically test our metrics for their validity and reliability. Finally, we explore and find a relationship between the scores and the model’s expected familiarity with an entity, and provide two use cases to illustrate their benefits.
@@ 1 Introduction
Language models have displayed remarkable abilities to answer factual queries about entities, suggesting that they encode knowledge about these entities learned during pretraining (Petroni et al., 2019; Brown et al., 2020; Roberts et al., 2020; Geva et al., 2021). For prompts that extend a question with additional information or context, the model can draw on both its prior knowledge and the additional context to answer the query (Kwiatkowski et al., 2019; Joshi et al., 2017; Berant et al., 2013; Kasai et al., 2023). While previous research has investigated how often a model will rely on prior knowledge over conflicting contextual information in answering questions (Longpre et al., 2021), we hypothesize that models will not behave identically for all contexts and entities. For example, if a language model is prompted with Harry hugged Voldemort. How friendly are Harry Potter and Lord Voldemort?, we might expect the prior knowledge learned from training data describing the rivalry between these two characters to significantly influence the model’s answer. However, if the model lacks a strong prior on, say, Susie and Alia, then we might expect its answer to be primarily context-driven when prompted with Susie hugged Alia. How friendly are Susie and Alia?.
@@
We formalize this problem through the lens of evaluating the change in a model’s answer distribution for different contexts and entities. We present two mutual information-based metrics that allow us to explore differences in the effect of specific contexts on model behavior for different entities. The persuasion score of a given context measures how much a model’s answer distribution is affected by the context when prompted with a particular query about a given entity. The susceptibility score of a given entity measures how much the model’s answer distribution can be swayed for a particular query about that entity, marginalized over all contexts. Given their basis in mutual information, our metrics are natural operationalizations of persuasion and susceptibility. Furthermore, we offer empirical evidence of the validity and reliability of these measures by comparing them against similar metrics and showing their robustness to paraphrases and different samples.
@@
To study how language models behave for different contexts and entities, we create a synthetic dataset of queries covering 122 topic areas extracted from the YAGO knowledge graph (Suchanek et al., 2007), entities extracted from YAGO and generated with GPT-4 OpenAI (2023), and contexts constructed with different qualities, e.g., relevancy and assertiveness. We apply our new metrics to Pythia models ranging from 70m to 12b parameters (Biderman et al., 2023) to find evidence that relevant contexts are consistently more persuasive than irrelevant ones, and assertive contexts are more persuasive than less assertive ones for yes–no questions. In a deep dive into one model, we find evidence that entities with high expected familiarity, as measured by both training data frequency and entity degree statistics in the YAGO knowledge graph, have lower susceptibility scores.
@@
We further conduct case studies to show how these metrics could be useful in applied settings. In a study on friend–enemy stance measurement, we find evidence that enemy duos are less susceptible than friend duos. Applying our metrics to gender bias, we find evidence for a difference in susceptibility between stereotypically masculine and feminine names for gender-biased contexts. Through this, we show how our proposed metrics can be used to better analyze the effects of context and prior knowledge, with the potential for application toward greater control over model behavior.
@@ 2.1 Entity Bias
Loosely defined across studies as model bias from spurious correlations between entity mentions and a target characteristic in the training data, entity bias has been mainly examined in relation extraction (RE), where a model extracts relationships between entities from text (Zelenko et al., 2002; Zhou et al., 2005). Different studies on RE have attempted to either mitigate (Zhang et al., 2017, 2018; Peng et al., 2020; Wang et al., 2022) or leverage (Yamada et al., 2020; Zhou and Chen, 2022) entity bias for improved performance on the task. Wang et al. (2022) note that entities can carry both useful semantic information about their roles, e.g., whether the entity is a person or a place, and spurious information which can bias the model toward relations not mentioned in the target sentence, e.g., the model may link Switzerland with the relation countries_of_residence at inference time for a sentence that does not mention this relation, due to their frequent co-occurrence in the training data. Additional studies on machine reading comprehension (MRC) tasks, e.g., SQuAD (Rajpurkar et al., 2016), TriviaQA (Joshi et al., 2017), and NaturalQuestions (Kwiatkowski et al., 2019), have found substituting different entity names can result in meaningful changes in model predictions and overall evaluation performance (Yan et al., 2022).
@@ 2.2 Context and Entity Bias
The existence of entity bias naturally raises the question of how it interacts with context to shape a model’s response. Several papers (Longpre et al., 2021; Chen et al., 2022; Xie et al., 2023) approach this by inducing and exploring knowledge conflicts, where a context preceding a query proposes information that conflicts with a model’s prior knowledge about that query. They measure the model’s reliance on pretrained entity bias and context by computing the memorization ratio: the proportion of knowledge conflict examples for which the model maintains its answer from prior knowledge. Pezeshkpour (2023) measures a model’s prior knowledge of a fact by comparing the entropy of a queried model’s answer distribution before and after stating the fact in context. Several studies further propose interventions to reduce entity bias and favor in-context information. These approaches include prompting (Zhou et al., 2023; Onoe et al., 2023), modifying training data (Wang et al., 2023), fine-tuning (Li et al., 2023), and neuron-level interventions (Yu et al., 2023) at inference time.
However, the metrics used to quantify model reliance on context and entity bias in these papers—excepting Pezeshkpour (2023)—are limited in several ways. First, most previous work does not develop entity-specific or context-specific metrics for the strength of the entity bias or context persuasiveness. Instead, their metrics produce only a single number to summarize the model’s overall reliance on entity bias. Second, their setups are limited to adversarial cases in which a context is chosen to go against the entity bias. Indeed, we may well wish to measure the interplay of context and entity bias in other cases, such as when context reinforces entity bias or when they do not clearly disagree. Therefore, we seek to define a metric that measures how much the model depends on a given context or entity with rigorous, theoretically grounded interpretations.
@@ 3 Our Formalization
We now formalize the problem setting in which we define metrics for the susceptibility of an entity and the persuasiveness of a context for a given model. Let be an alphabet. Consider a language model over , i.e., is a distribution over the Kleene closure . Furthermore, we assume that was estimated from a corpus . Let be the subset of strings that could correspond to string representations of entities.111In practice, an entity can have different verbalizations, e.g., Sherlock and Sherlock Holmes. . And, whether a specific string refers to an entity may very well depend on the context in which it occurs. Thus, our set is an approximation at best. Let be a set of slotted query templates of type that fill the slot with the argument to each . We can think of a query template as slotting an entity into a query, e.g., slotting the entity Slovenia into the query template The capital of is produces the string The capital of Slovenia is.
Now, consider three random variables. First, let be a -valued random variable that stands for a context. Second, let be a -valued random variable whose values are answers. Let be a -valued random variable. The pushforward , then, is a -valued random variable over slottings of entities according to query . The random variables , and are jointly distributed according to the following probability distribution
where is string concatenation. To formalize persuasion and susceptibility, we make use of the joint distribution heavily in the proceeding subsections.
@@ 3.1 Persuasion Score
For each context that is prepended to a query template with a given entity slotted in, , we wish to assign a persuasion score to represent how successful that context is at altering a model’s answer distribution. This score depends on the specific queried entity , because contexts themselves are often entity-dependent. Intuitively, a context’s persuasion score should measure how much the probability distribution of possible answers changes, averaged across all possible answers. More precisely, we define our persuasion score as the half-pointwise mutual information (half-PMI) between the context and the answer random variable , conditioned on the fixed query about an entity:
where and can be derived by marginalizing and conditioning Eq. 1.
The persuasion score of a context can then be interpreted as the degree (in nats) to which the context was able to change the model’s answer distribution when prepended to a query. When the persuasion score is at its lower bound of 0 nats, it indicates the context is completely unpersuasive, i.e., it did not change the model’s answer distribution at all. Contexts with higher persuasion scores change the answer distribution more, which is consistent when viewed through the lens of KL-divergence.222We include details on further equivalencies of half-PMI and other concepts in information theory in App. A.
@@ 3.2 Susceptibility Score
For a given query template applied to an entity , we further wish to assign a susceptibility score to which represents how easy it is to change a model’s answer distribution. Intuitively, the susceptibility score should measure how much a model’s answer distribution to a query changes when prompted with additional context, averaged across all possible contexts and answers. More precisely, we define the susceptibility score as the mutual information between the context and answer random variables, conditioned on a fixed query about an entity:
where is the marginal distribution over contexts. Equivalent to the difference in entropy , the susceptibility score represents the reduction in answer distribution uncertainty (in nats) when a query is preceded by a context. A high susceptibility score means the model is highly influenced by context for the query about that entity, with its upper bound of indicating that context fully determines the answer. A low score indicates the model’s response is robust to context, with its lower bound of indicating no influence of context on the answer distribution. We can use susceptibility to answer the question “How much does the model’s answer depend on its entity bias?” with an information-theoretically grounded, interpretable scale based on the model’s full behavior.
@@ 4.1 Setup
For each of the 122 different relations from the YAGO knowledge graph, we collect 100 entities33350 are real entities (e.g., Adele) sampled from YAGO, and 50 are fake entities (e.g., Udo König) of the same entity class (e.g., Person) generated with GPT-4 (OpenAI, 2023). and construct 600 random contexts from relation-specific context templates such that each entity is mentioned in 6 contexts. The 600 contexts are evenly distributed between assertive,444E.g., Definitely, the capital of {entity} is {answer}. base,555E.g., The capital of {entity} is {answer}. and negation666E.g., The capital of {entity} is not {answer}. context types. We construct four query forms of each relation: two closed questions, i.e., yes–no, and two open questions.
Using these samples, we compute persuasion scores for each context according to Eq. 2 and susceptibility scores for each entity according to Eq. 3 for each of the four query forms for six Pythia models of different sizes77770m, 410m, 1.4b, 2.8b, 6.9b, 12b (8-bit quantized) trained on the deduplicated Pile (Wolf et al., 2020; Dettmers et al., 2022; Biderman et al., 2023). Due to computational constraints, we approximate the model’s answer distribution with the next-token distribution over the model’s vocabulary.888While it would be more precise to estimate the answer distribution by repeatedly sampling many model outputs, this is very computationally expensive. The detailed setup can be found in App. C.
@@ 4.2 Empirically Validating Our Metrics
4.2.1 Estimating Scores
Because persuasion and susceptibility scores both involve a countably infinite sum, we opt for a stochastic approximation scheme. Specifically, we construct a Monte Carlo estimator. We sample from a narrower set of constructed contexts and approximate the answer distribution with the next token distribution over the model’s vocabulary, as described in § 4.1. We take a sample size of 600 contexts. While the Monte Carlo approximation itself results in a consistent estimator, the additional approximations mean we do not have a guarantee on the quality of the approximation as a whole. In Fig. 4, we exhibit the variance of our estimator across three random seeds, i.e., sampled sets of context.
@@ 4.2.2 Validating Persuasion Scores
According to existing measurement modeling methods (Loevinger, 1957; Messick, 1987; Jackman, 2008; Hand, 2004; Quinn et al., 2010; Jacobs and Wallach, 2021), observing a relationship between a new metric and existing ones would serve as additional evidence that the metric is meaningful. To this end, we explore whether contexts with higher persuasion scores tend to more successfully convince the model to agree with the context. Using the Pythia-6.9b-deduped model, we generate an answer for each prepended context to a query–entity pair from § 4.1 and use simple string matching to map the answer to whether it agrees with the context, the original answer, or neither. We then apply a permutation test ( with the Benjamini–Hochberg (BH) correction (Benjamini and Hochberg, 1995)) for whether contexts that elicited context-concordant answers have higher persuasion scores than those that did not alter the model from the original answer for each of the 122 query topics. Within a query topic, we run separate tests for open and closed queries, since the entropy of the answer distribution for closed queries tends to be much lower than the entropy for open queries. We find that for of open queries, contexts that persuade the model to output the in-context answer have significantly higher persuasion scores than non-persuasive ones. Curiously, this behavior holds for only of closed queries. We discuss this surprising result more in § 6. We further show a summary of the mean persuasion scores for the different kinds of elicited answers in Fig. 2.